OpenStudy (ilovemath14):

HELP ASAP!! DUE TOMORROW!! FAN AND MEDAL!! 4. $4000 principal earning 4% compounded annually, 5 years The balance after 5 years will be $--------- 6. 3,700 deposit earning 3.3% compounded monthly, after 5 years. The balance after 5 years will be $--------

OpenStudy (4everaddicted2anime):

The formula for this is \[A(t)=P(1+\frac{ r }{ n })^{nt}\]

OpenStudy (4everaddicted2anime):

r= rate n= number of times compounded t= time p= principle

OpenStudy (ilovemath14):

A(5)= 4000(1+0.04)^5 @4everaddicted2anime

OpenStudy (anonymous):

Multiply the $4000 by .04 and multiply by 5 I think

OpenStudy (ilovemath14):

@myininaya @Michele_Laino @johnweldon1993 @KendrickLamar2014 @jabez177 @Preetha

OpenStudy (michele_laino):

I'm sorry, I'm not good with financial mathematics

OpenStudy (anonymous):

A(5)= 4000(1+0.04)^5 = 4000*(1.04^5) = ? Just plug into a calculator.